OpenStudy (anonymous):

how do i solve the equation cos(x)+2cos(x)sin(x)=0 on the interval [0,2π)?

OpenStudy (anonymous):

Pull out a common factor: cosx + 2 cosx sinx = 0 cosx (1 + 2 sinx) = 0 The zero product property gives you two equations: cosx = 0 1 + 2 sinx = 0, or sinx = -1/2 Solve for x in the given interval.

OpenStudy (amoodarya):

also this is a solution \[\cos(x)+2\cos(x)\sin(x)=0 \\cos(x)=-2\cos(x)\sin(x)\\cos(x)=-\sin(2x)\\cos(x)=\sin(-2x)\\cos(x)=\cos(\frac{ \pi }{ 2 }-(-2x))\\cos(x)=\cos(\frac{ \pi }{ 2 }+2x)\\x=\pm (\frac{ \pi }{ 2 }+2x)+ 2\pi k\]

OpenStudy (anonymous):

thanks!!!